# On the future cover of a sofic shift

### v1

@@ 4.2. Subset constructions for sofic shifts
commutes. When we use the same symbol $\kappa$ for the two conjugacies $\kappa:X_{G}^{L_{G}}\to X_{G^{\prime\prime}}$ and $\kappa:X_{H}^{L_{H}}\to X_{H^{\prime\prime}}$ from Lemma 4.2, we note that $\phi^{\prime\prime}=\kappa\circ\phi^{\#}\circ\kappa^{-1}$. Since $\phi^{\#}(L_{G}^{-1}(y))=L_{H}^{-1}(\psi(y))$ it follows that $\phi^{\prime\prime}(\beta_{G}(y))=\beta_{H}(\psi(y))$ for all $y\in Y$. By Lemma 4.7 this implies that $\phi^{\prime\prime}(X_{G^{\prime}})=X_{H^{\prime}}$, and we set $\phi^{\prime}:=\phi^{\prime\prime}|_{X_{G^{\prime}}}$. This proves the existence part. To prove uniqueness, consider a conjugacy $\lambda:X_{G^{\prime}}\to X_{H^{\prime}}$ such that
@@
${X_{G}^{L_{G}}\cap\kappa^{-1}(X_{G^{\prime}})\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces}$$\scriptstyle{\lambda^{\prime}}$$\scriptstyle{L_{G}^{\#}}$${X_{H}^{L_{H}}\cap\kappa^{-1}(X_{H^{\prime}})\ignorespaces\ignorespaces\ignorespaces\ignorespaces}$$\scriptstyle{L_{H}^{\#}}$${Y\ignorespaces\ignorespaces\ignorespaces\ignorespaces}$$\scriptstyle{\psi}$${Z}$
commutes. Note that $L_{G}^{-1}(y)\in X_{G}^{L_{G}}\cap\kappa^{-1}(X_{G^{\prime}})$ by Lemma 4.7. It follows that $\lambda^{\prime}(L_{G}^{-1}(y))\subseteq L_{H}^{-1}(\psi(y))$ and ${\lambda^{\prime}}^{-1}(L_{H}^{-1}(\psi(y)))\subseteq L_{G}^{-1}(y)$, implying that $\lambda^{\prime}(L_{G}^{-1}(y))=L_{H}^{-1}(\psi(y))$ for all $y\in Y$ and hence $\lambda(\beta_{G}(y))=\beta_{H}(\psi(y))=\phi^{\prime}(\beta_{G}(y))$ for all $y\in Y$. By Lemma 4.7 this implies that $\lambda=\phi^{\prime}$.
@@ 5.2. A cover of the future cover
The start vertex $s_{\overline{G}}(D,a)$ of $(D,a)\in E_{\overline{G}}$ is $D\in 2^{V_{G}}$ and the terminal vertex $t_{\overline{G}}(D,a)$ is $[D,a]$. With this definition, there is a graph morphism $G\to\overline{G}$ sending a vertex $v\in V_{G}$ to $\{v\}\in 2^{V_{G}}$ and $e\in E_{G}$ to $(\{s_{G}(e)\},L_{G}(e))\in E_{\overline{G}}$, and we extend $L_{G}$ from $E_{G}$ to a labeling $L_{\overline{G}}$ of $E_{\overline{G}}$ such that
@@ 6. Four examples
This following labeled graph $(G,L_{G})$ is irreducible, right-resolving and follower-separated and hence the minimal right resolving graph of the sofic shift $Y$ it represents. Note that $Y$ is mixing.

### v2

@@ 4.2. Subset constructions for sofic shifts
commutes. When we use the same symbol $\kappa$ for the two conjugacies $\kappa:X_{G}^{L_{G}}\to X_{G^{\prime\prime}}$ and $\kappa:X_{H}^{L_{H}}\to X_{H^{\prime\prime}}$ from Lemma 4.2, we note that $\phi^{\prime\prime}=\kappa\circ\phi^{\#}\circ\kappa^{-1}$. Since $\phi^{\#}(L_{G}^{-1}(y))=L_{H}^{-1}(\psi(y))$ it follows that
(14) $\phi^{\prime\prime}(\beta_{G}(y))=\beta_{H}(\psi(y))$
for all $y\in Y$. Let $\mathcal{C}_{G}$ denote the set of components $C$ of $G^{\prime\prime}$ with the property that there is a point $y\in Y$ such that $\beta_{G}(y)$ is backward asymptotic to $X_{C}$. We denote the analogous set of components in $H^{\prime\prime}$ by $\mathcal{C}_{H}$. It follows from (14) that when $C\in\mathcal{C}_{G}$ we have that $\phi^{\prime\prime}(X_{C})=X_{C^{\prime}}$ for some $C^{\prime}\in\mathcal{C}_{H}$. If $x\in X_{G^{\prime\prime}}$ is backward asymptotic to $X_{C}$ we have that $\phi^{\prime\prime}(x)$ is backward asymptotic to $X_{C^{\prime}}$. Since the elements of $X_{G^{\prime\prime}}$ that are backward asymptotic to $X_{C}$ for some $C\in\mathcal{C}_{G}$ constitute a dense set in $X_{G^{\prime}}$, it follows that $\phi^{\prime\prime}(X_{G^{\prime}})\subseteq X_{H^{\prime}}$. By symmetry we have equality; $\phi^{\prime\prime}(X_{G^{\prime}})=X_{H^{\prime}}$, and we set $\phi^{\prime}:=\phi^{\prime\prime}|_{X_{G^{\prime}}}$. This proves the existence part. To prove uniqueness, consider a conjugacy $\lambda:X_{G^{\prime}}\to X_{H^{\prime}}$ such that
@@
(15) ${X_{G}^{L_{G}}\cap\kappa^{-1}(X_{G^{\prime}})\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces}$$\scriptstyle{\lambda^{\prime}}$$\scriptstyle{L_{G}^{\#}}$${X_{H}^{L_{H}}\cap\kappa^{-1}(X_{H^{\prime}})\ignorespaces\ignorespaces\ignorespaces\ignorespaces}$$\scriptstyle{L_{H}^{\#}}$${Y\ignorespaces\ignorespaces\ignorespaces\ignorespaces}$$\scriptstyle{\psi}$${Z}$
commutes. Note that $L_{G}^{-1}(y)\in X_{G}^{L_{G}}\cap\kappa^{-1}(X_{G^{\prime}})$ for all $y\in Y$ since $\beta_{G}(Y)\subseteq X_{G^{\prime}}$ by definition. It follows from (15) that $\lambda^{\prime}(L_{G}^{-1}(y))\subseteq L_{H}^{-1}(\psi(y))$ and by symmetry that $\lambda^{\prime}(L_{G}^{-1}(y))=L_{H}^{-1}(\psi(y))$ for all $y\in Y$. Hence $\lambda(\beta_{G}(y))=\beta_{H}(\psi(y))$ for all $y\in Y$. In particular, $\lambda(\beta_{G}(y))=\phi^{\prime}(\beta_{G}(y))$ for all $y\in Y$ and hence, by Lemma 4.4, $\lambda|_{X_{C}}=\phi^{\prime}|_{X_{C}}$ for $C\in\mathcal{C}_{G}$. Consider then an $x\in X_{G^{\prime}}$ backward asymptotic to $X_{C},\ C\in\mathcal{C}_{G}$. Since $\lambda$ and $\phi^{\prime}$ agree on $X_{C}$ it follows that there is an $N\in\mathbb{Z}$ such that $\lambda(x)_{(-\infty,N]}=\phi^{\prime}(x)_{(-\infty,N]}$. Since $L_{H^{\prime}}$ is right-resolving and $L_{H^{\prime}}(\lambda(x))=\psi(L_{G^{\prime}}(x))=L_{H^{\prime}}(\phi^{\prime}(x))$, we find that $\lambda(x)=\phi^{\prime}(x)$. Since such $x$’s are dense in $X_{G^{\prime}}$ we conclude that $\lambda=\phi^{\prime}$. ∎
@@ 5.2. A cover of the future cover
The start vertex $s_{\overline{G}}(D,a)$ of $(D,a)\in E_{\overline{G}}$ is $D\in 2^{V_{G}}$ and the terminal vertex $t_{\overline{G}}(D,a)$ is $[D,a]$. We define a labeling $L_{\overline{G}}$ of $\overline{G}$ such that
@@ 6. Three examples
The following labeled graph $(G,L_{G})$ is irreducible, right-resolving and follower-separated and hence the minimal right resolving graph of the sofic shift $Y$ it represents. Note that $Y$ is mixing.